### v1

# A Survey on LLM Test-Time Compute via Search: Tasks, LLM Profiling, Search Algorithms, and Relevant Frameworks

Abstract
LLM test-time compute (or LLM inference) via search has emerged as a promising research area with rapid developments. However, current frameworks often adopt distinct perspectives on three key aspects—task definition, LLM profiling, and search procedures—making direct comparisons challenging. Moreover, the search algorithms employed often diverge from standard implementations, and their specific characteristics are not thoroughly specified. In this survey, we provide a comprehensive technical review that unifies task definitions and provides modular definitions of LLM profiling and search procedures. The definitions enable precise comparisons of various LLM inference frameworks while highlighting their departures from conventional search algorithms. We also discuss the applicability, performance, and efficiency of these methods. For further details and ongoing updates, please refer to our GitHub repository: https://github.com/xinzhel/LLM-Agent-Survey/blob/main/search.md.
1 Introduction
Scaling test-time compute via search has recently enhanced the LLMs’ power to a new level on reasoning tasks (Yao et al., 2023a; Hao et al., 2023), sequential decision-making tasks (e.g., robotics) (Putta et al., 2024), and graph-traversal tasks (e.g., path finding) (Meng et al., 2024). This survey aims to provide a comprehensive but integrated survey on existing frameworks for LLM-based search. The focus is on the work that the search processes are coupled with LLMs’ test time compute rather than those using search and LLMs separately. For example, the plans (commonly in the form of PDDL) are prepared by LLMs to perform local search (Valmeekam et al., 2023b; Guan et al., 2023; Valmeekam et al., 2023a).
1.1 Existing Surveys
Current reviews on LLM search are limited from the following perspectives.
No dedicated, Detailed Survey
Current surveys only contain paragraphs or sections to roughly touch on both technical aspects and their practical applicability, as summarized in Table 1.
Limited Mention on LLM-Side Design
Specifically, most of existing surveys (Huang et al., 2024; Wang et al., 2024b) mention little or a few implementations and dimensions for LLM profiling, which is not suitable for all the frameworks. Besides, the lack of examples hinders understanding.
Limited Mention on Search
Li (2024) give more detail regarding LLM profiling but lack details on search processes. Nonthelessness, most of existing surveys (Huang et al., 2024; Wang et al., 2024b) give a general sense of the computation process by mentioning which classical search algorithms the frameworks are built upon (e.g., depth-first search). However, details should be given because of their nuanced differences. Besides, many untypical twists to classic search algorithms are hidden. The deviations are not friendly for newcomers in the newly-developed area, e.g., those computer science graduates educated with typical search algorithms.
Table 1: Comparisons with other surveys related to LLM-integrated search. “Coverage” indicates the number of related papers, “Sections” refers to the sections of the survey manuscripts that discuss LLM inference via search, and “Def.” is the abbreviation for Definition. The URLs for the reference versions are given at the footnotes. “Mentioned” refers to the mention of the LLM function, while “Limited” means that a formal definition is given without specific distinctions. The numerical values correspond to the number of implementations (impl.) or dimensions (dim.) identified for LLM-Profiled Roles (LMPRs); different dimensions may lead to a combinatorial number of implementations. The “11” in “11+” specifically refers to the frameworks we describe in detail.
Coverage Sections Task Def. Search Algorithms LLM Profiling
Details Deviations Policy Value Transition
Huang et al. (2024) 5 ✓(§4) 1 impl. 1 impl.
Wang et al. (2024b) 3 ✓(§2.1.3) Mentioned
Li (2024) 8 ✓(§4.3) Limited 2 impl. 2 dim. NA
Ours 11+ All 4 impl. 4 dim.; 11 impl. 2 impl.
1.2 Survey Structure
To solve the above limitations, we provide unified task definitions and decouple the LLM-specific design (mainly prompting) from the control program (search procedures/algorithms). There exists a hierarchical structure between them: the low-level definitions provide a unified interface for the high-level components. The overall structure, accompanied by illustrative examples, is presented in Figure 1.
Introducing a Unified Task Definition Based on MDPs
(§ 2) Our definition standardizes different tasks in MDP structure. While MDPs naturally align with AI domains like robotics, special attention is given to adapting this definition for tasks traditionally not modeled as MDPs, such as graph traversal, reasoning, dialogue systems, and code generation. Notably, this MDP-based definition is also applicable to other LLM inference frameworks beyond search, including works like Li et al. (2022), Zhao et al. (2023), and Hu et al. (2024).
Comprehensively Summarizing LLM Profiling and Implementations
(§ 3) The design and implementation of LLM profiling and prompting can be modularized into components commonly used in solving MDPs (Sutton & Barto, 2018): policies, value functions, and transition models. Correspondingly, 3 types of LLM-Profiled Roles (LMPRs) are defined.
Defining Modular Search Procedures
(§ 4) Rather than directly showcasing individual search-based frameworks for LLM inference, we focus on modular and reusable components to reduce redundancy and enable more straightforward comparisons across frameworks. This approach promotes flexibility and minimizes overhead when adapting or extending search methods.
Reviewing Individual Frameworks
(§ 5) Based on the unified task and LMPR interface, we provide a comprehensive review of individual frameworks, organized by the search algorithms they are built upon. Our analysis highlights how LLM integration either diverges from or enhances traditional search algorithms. We identify and clearly present 11 frameworks, summarized in Table 7. This count exclusively includes frameworks that focus on test-time computation through search detailed in Section 5. Additionally, we highlight other test-time frameworks that function as components within search processes, such as ReAct (Yao et al., 2023b), CoT (Wei et al., 2022), and Self-Consistency (Wang et al., 2023), along with those discussed in Section 7.
Analyzing Key Perspectives of LLM-Integrated Search Methods
(§ 6) We critically examine these methods from four key perspectives: deviation, applicability, performance, and efficiency. For deviations, we compare the structural and functional differences between LLM-integrated frameworks and standard search algorithms as described in foundational texts such as Russell & Norvig (2010) and Sutton & Barto (2018). This analysis highlights how LLMs modify or enhance traditional search processes, providing a deeper understanding of their impact and potential.
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7em,font=,, where level=2text width=8.5em,font=,, where level=3text width=10.5em,font=,, where level=4text width=12em,font=,, [ LLM Search, ver [ Task
Definitions  (§2) [ Web
Navigation (§2.1) [ E.g.,  Agent Q (Putta et al., 2024), Koh et al. (2024) , LATS (Zhou et al., 2024a) , leaf, text width=44.5em ] ] [ Graph
Traversal (§2.2) [ E.g.,  LLM-A* (Meng et al., 2024) , leaf, text width=44.5em ] ] [ Reasoning (§2.3) [ via Concatenation [ E.g.,  Li (2024), Q* (Wang et al., 2024a), CPO (Zhang et al., 2024) , leaf, text width=32.5em ] ] [ via LLMs [ E.g.,  RAP (Hao et al., 2023), LeastToMost  (Zhou et al., 2023) , leaf, text width=32.5em ] ] [ via Concatenation
and Tool Invocation [ E.g.,  LATS (Zhou et al., 2024a), ReAct (Yao et al., 2023b) , leaf, text width=32.5em ] ] ] [ Code
Generation (§2.4) [ E.g.,  PG-TD (Zhang et al., 2023), LATS (Zhou et al., 2024a) , leaf, text width=44.5em ] ] [ Goal-oriented
Dialog (§2.5) [ E.g.,  GDP-ZERO (Yu et al., 2023) , leaf, text width=44.5em ] ] ] [ Low-level
Modules:
LLM-
Profiled
Roles (§3) [ Policy (§3.1) [ Naive Policy [ E.g.,  Xie et al. (2023) , leaf, text width=32.5em ] ] [ ReAct Policy [ E.g.,  ReAct (Yao et al., 2023b) , leaf, text width=32.5em ] ] [ Batch Policy [ E.g.,ToT(Yao et al., 2023a) , leaf, text width=32.5em ] ] ] [ Transition
Model (§3.2) [ E.g.,RAP(Hao et al., 2023) , leaf, text width=32.5em ] ] [ Evaluator (§3.3) [ E.g.,RAP(Hao et al., 2023), Xie et al. (2023) , leaf, text width=32.5em ] ] ] [ Mid-level
Modules:
Search
Procedures
4) ] [ Frameworks
Based on
High-level
Search
Algorithms (§5) [ Beam Search [ E.g., Xie et al. (2023),PathFinder (Golovneva et al., 2023) , leaf, text width=44.6em ] ] [ BFS (breath) [ E.g., ToT(Yao et al., 2023a) , leaf, text width=44.6em ] ] [ DFS [ E.g., ToT(Yao et al., 2023a) , leaf, text width=44.6em ] ] [ BFS (Best) [ E.g., Koh et al. (2024) , leaf, text width=44.6em ] ] [ A* [ E.g., LLM-A* (Meng et al., 2024),Q* (Wang et al., 2024a) , leaf, text width=44.6em ] ] [ MCTS [ E.g.,  RAP (Hao et al., 2023), LATS (Zhou et al., 2024a), (Zhao et al., 2023) , leaf, text width=44.6em ] ] ] [ Related
Work
7) [ LLM fine-tuning
for LMPRs [ E.g.,  CPO (Zhang et al., 2024), AlphaZero-like tree search (Wan et al., 2024) , leaf, text width=44.6em ] ] [ Search with
Multi-Modal LLMs [ E.g.,  Mulberry(Yao et al., 2024) , leaf, text width=44.6em ] ] [ Branching without
Search [ E.g.,  Tree-Planner (Hu et al., 2024), Graph-of-Thoughts (GoT) (Besta et al., 2024) , leaf, text width=44.6em ] ] ] ]
Figure 1: Survey structure. This table shows a selected subset of the works reviewed. Related work concerning search procedures is not included here, as these procedures are largely consistent across different frameworks and listing them in the graph would not aid in differentiating the works. Comprehensive lists of related work are provided in the following tables: Table 2 for task definitions, Table 4 for LLM-Profiled Evaluator, and Table 7 for LMPRs and search procedures.
1.3 Intended Audience and Use Cases
Reusable Modules
While we strive to provide comprehensive coverage of the latest research, we acknowledge the rapid pace of advancements in the field, where variations in LLM profiling and search implementations may not be covered. Nonetheless, this survey offers a collection of classical and reusable implementations that can serve as solid foundations for future research and development.
Anchoring Purpose
Our work serves as a valuable reference in two key ways: 1) Incorporating Novel Designs with Minimal Adjustments: During the development of this survey, we seamlessly integrated emerging methods with minor modifications. For example, we separated single-step simulation from path simulation in MCTS to accommodate LATS (Zhou et al., 2024a) and decoupled value-based selection from LMPE+ or state-based evaluation to allow specialized ways to assign values, e.g., using parents’ states for evaluation Koh et al. (2024). 2) Expanding with Additional Details: Although we intentionally omit details unrelated to search procedures—such as how the reflection module is incorporated in LATS (Zhou et al., 2024a)—our structured presentation of search control flows serves as a stable anchor for integrating such complex components. Although many of these design elements are non-trivial, our framework simplifies readers’ understanding of their integration by using the search process as a guiding anchor.
For Research Engineers
To support practical implementation, search procedures (Section 4) are presented in an Object-Oriented Programming (OOP) style, promoting modularity and ease of integration for research engineers.
Limitations
We adopt Markov Decision Process (MDP) definitions to unify and compare various methods due to their comprehensive nature. However, this formalism may feel excessive for readers focused on specific frameworks or tasks where transition and action definitions are unnecessary. For instance, the Tree-of-Thoughts (Yao et al., 2023a) approach could be more intuitively understood without relying on MDP foundations.
Table 2: A unification of various tasks for LLM-based search.
Tasks Actions States / Observations Transitions Rewards Action Reversible? Related LIS Works
Web Navigations Discrete,
constrained,
heterogeneous
Discrete,
infinite
Dynamic or Deterministic 1 if the goal is achieved, e.g., buying a coffee mug Maybe Agent Q (Putta et al., 2024); Koh et al. (2024)
Graph traversal Discrete,
constrained,
homogeneous
Discrete,
finite (commonly)
Deterministic 1 if $s\in G$ Maybe LLM-A* (Meng et al., 2024)
Reasoning ($\mathcal{T}$ via Concatenation) Open (A thought expressed in one or more tokens) Open, unknown until reached (Problem description + Concatenated thoughts) Deterministic (Concatenating) 1 if the final results $=$ ground-truth Q* Wang et al. (2024a) CPO (Zhang et al., 2024); ToT (Yao et al., 2023a); RStar (Anonymous, 2024); MindStar (Kang et al., 2024)
Reasoning via QAs Open Open, unknown until reached Dynamic (Question answering then concatenating Q&A) 1 if the final results $=$ ground-truth RAP (Hao et al., 2023); RStar (Anonymous, 2024) ; LeastToMost (Zhou et al., 2023);
Reasoning ($\mathcal{T}$ via concatenations and tool invocation) Open Open, unknown until reached Deterministic (Concatenating); dynamic or deterministic (tool invocation) 1 if the final results $=$ ground-truth Maybe LATS (Zhou et al., 2024a)
Code Generation Open (A single token) Open, unknown until reached (Problem description + Concatenated tokens) Deterministic pass rate of the complete program PG-TD (Zhang et al., 2023)
Goal-oriented Dialog Discrete, constrained (An intent) A sequence of intents, agent/user utterances Dynamic 1 if the conversational goal is achieved GDP-Zero (Yu et al., 2023)
2 Task (Re)formulation
Tasks solved by LLM-integrated search are inherited from both the “LLM” side (human language tasks) and the “search” side (structured tasks): 1) language reasoning: LLMs are naturally applied to reasoning tasks in language (Wei et al., 2022). 2) structured tasks: On the other hand, search algorithms are more conventionally utilized for structured tasks, such as web navigation, robotic navigation, gaming, and graph traversal (Russell & Norvig, 2010; Sutton & Barto, 2018). The convergent nature is that all of them belongs to sequential decision-making, e.g., reasoning often involves generating and evaluating sequences of logical steps or decisions to arrive at a conclusion.
A MDP-Like Formulation
To enable a clear comparison across different frameworks, this section formulates the tasks in Markov Decision Processes (MDPs) $\langle S,A,\mathcal{T},R\rangle$. In addition, observations $O$ are often considered in Partially Observable Markov Decision Processes (POMDPs) (Li et al., 2022; Zhao et al., 2023; Hu et al., 2024):
  • A set of states $S$, including the goal state $s_{g}$;
  • A set of observations $O$, where $o_{t}$ is the partial observations of the state $s_{t}$ at time step $t$;
  • A set of actions $A$, where $a_{t}\in A$ is the action on $s_{t}$;
  • Transitions $\mathcal{T}\left(s_{t+1}\mid s_{t},a_{t}\right)$, which define the dynamics from one state to another after executing an action;
  • Rewards $R$: A rewards evaluate the "quality" of a state-action pair or a trajectory towards the desired outcomes or goals.
This section only discuss task elements that are external to agents and exist independently of how agents operate or learn. We will see in the next section how the POMDP setting fits in LMPRs and agent definitions, where the Markovian assumption is broke.
A Summary of Concrete Tasks
Some structured tasks (e.g., recycling robot, gridworld, and chess) are always modeled as MDPs. These typical MDP tasks are actively studied in the domains of reinforcement learning (Sutton & Barto, 2018), intelligent agents, and control theory. Actions and states are always explicitly defined. Commonly, a physical environment or a rule set defines a discrete and finite action space. And states are commonly finite and can be enumerated, e.g., all the possible configurations of chess board or the grid areas robot can travel. However, others, e.g., graph traversal and reasoning tasks, are not commonly formalized as MDPs until the emergent of LLM-based agents. In particular, the MDP elements in graph traversal and reasoning tasks are not explicit. The rest of this section mainly discusses how the these tasks fit the following MDP notations, as summarized in Table 2.
2.1 Web Navigation
Another type of tasks is to navigate on websites for shopping and retrieving information (Zhou et al., 2024b)
  • States/observations: A state/observation is normally a web page. For example, the beginning state can be the homepage. The transition is governed by a deterministic transition function. The state is not always accessible to the agent. For example, “$s_{t}$ may include private information such as database entries of the site” (Koh et al., 2024)).
  • Reward: The reward is given when the goal $g$ is reached, e.g., successfully ordering a product.
2.2 Graph Traversal in MDPs
A graph traversal problem, e.g., robotic navigation, is represented as a graph $G=(V,E)$, where $V$ is a set of vertices (or nodes), and $E\subseteq V\times V$ represents the transitions between them. However, this definition is the common task settings for uninformed and informed search. However, these algorithms integrated with LLMs can be generalized beyond this definition. This is why recent work (Yao et al., 2023a) that uses these search algorithms along with LLM often uses MDP terminology, but does not include formal unification. Hence, we propose a conceptual framework that views graph traversal as a simplified version of MDPs. This framework can be described as follows:
  • States: Each node $v\in V$ is a state.
  • Actions: Actions are represented by edges $E$. The action space is generally considered homogeneous because the type of action is uniform, such as “MOVE[arg]”, where “[arg]” represents parameters like direction or target node.
  • Transitions: Following an edge from a node $s$ via an action (edge) $(s,s^{\prime})\in E$ always leads to the same node $s^{\prime}$. Hence, the transitions $\mathcal{T}$ are deterministic:
  • Rewards $R(s)$ are typically binary, with $R(s)=1$ if $s\in G$ (i.e., $s$ is a goal state) and $R(s)=0$ otherwise.
  • Heuristics for nodes $h(s)$ can be interpreted as estimates of the value function in the MDP formulation, representing an estimated cost-to-go from node $s$ to a goal node $g\in G$.
As best as we know, this conceptualization is not explicitly stated in any peer-reviewed literature.
2.3 Language Reasoning in MDP
The formulations of language reasoning tasks are more diverse and creative. Although not exhaustive, the following paragraphs summarize forms that are particularly used in the current study of LLM-integrated search.
Reasoning ($\mathcal{T}$ via Concatenation)
A reasoning process can be concretized as a Chain of Thoughts (CoT) $T_{1},T_{2},\ldots$ (Wei et al., 2022), each expressed as a sequence of tokens via LLM generation. The reasoning steps can be not only naturally evolved but also deliberately specified. For creative writing, the first step can be specified as planning, and the second step is to generate according to the plan (Yao et al., 2023b) Following previous work (Li, 2024; Wang et al., 2024a), the MDP formulation includes:
  • Actions: An action is a thought consisting of several tokens, i.e., $a_{1}=T_{1}$.
  • States: The initial state $s_{1}$ is defined by the task information, e.g., a user query, a problem description or the goal. The following states are defined as the concatenation of the following thoughts:
    $s_{t}=\left(s_{t-1},a_{t-1}\right)=\left(s_{1},a_{1},\ldots,a_{t-1}\right)$ (1)
    Apparently, directly concatnenating open actions leads to the open state space. When the reasoning is naturally evolved, the final state $s_{g}$ comes when the final thought $T_{g}$ or the entire chain expresses a valid response. It can be known in which step $s_{g}$ is reached for deliberate reasoning steps.
  • Transition $\mathcal{T}$: The deterministic state transition is defined for reasoning tasks. The next state $s_{t+1}$ is equal to the concatenation of $s_{t}$ and $a_{t}$.
  • Reward $R$: It is given when the final answer matches the ground truth and fits in human preference. This is normally integrated as the training objective of recent LLMs (Ouyang et al., 2022).
Reasoning via QAs
Some other works deliberately formulate an action space. A given task is decomposed into sequentially dependent subtasks (actions) requiring an Execute function, which relies on LLMs to solve. In other words, LLMs can be considered as transitions. Recent work on LLM search formulates a subtask as a question and use LLMs to answer the question. These subtasks (i.e., actions) can be generated all at once (Zhou et al., 2023) or in sequential order (Hao et al., 2023), each can be defined as an action $a_{t}$. $s_{t}$ is the concatenation of the task information $s_{0}$ and all the questions already answered with their answers: $\text{question}_{1},\text{answer}_{1},\ldots,\text{question}_{t},\text{answer}_{t}$.
Reasoning ($\mathcal{T}$ via Concatenation and Tool Invocations)
Once tool definitions are given to LLMs. During reasoning, LLMs can generate specific tokens to invoke tools. This integration of tool invocation into reasoning is firstly proposed by ReAct Yao et al. (2023b) and recently adapted to the LIS framework (Zhou et al., 2024a). Based on the definitions for Reasoning ($\mathcal{T}$ via Concatenation), the following things are added:
  • Actions: Tool-related actions are commonly discrete and constrained. For example, when Wikipedia is used, the possible actions include search[arg], lookup[arg]. Although not always, most works on LLM tool use only include the reading-only actions, the actions are reversible.
  • Transitions: Due to the change of the action space, the transitions can be either dynamic or deterministic, depending on the tools.
  • States: a state now concatenates not only the LLM generation but also tool responses. Moreover, LLM generation is not only about the direct thoughts for tasks but also the actions for tool invocations.
    $s_{t}=\left(s_{1},a_{1},o_{1},\ldots,a_{t-1},o_{t-1}\right)$ (2)
2.4 Code Generation in MDP
This is similar to language reasoning under Deterministic $\mathcal{T}$ via Concatenation. The only difference is that an action is a token in the vocabulary set of the LLM (rather than a thought consisting of several tokens). Such definition is originally proposed by Zhang et al. (2023). Under their definition, “the reward of state $s$ is the pass rate of the program on the public test cases. The reward of a partial program is always 0.”
2.5 Goal-Oriented Dialog in MDP
Previous work (Wang et al., 2020) frames goal-oriented dialog as MDP. Yu et al. (2023) begin using such formulation for LLM-integrated search. The formulation is demonstrated below.
  • Actions: An action $a\in A$ indicates the intent, which is predefined. For example, the intent to convince the Persuadee using reasoning and factual evidence is defined as “Logical Appeal”. This is commonly termed “dialog act” (Wang et al., 2020).
  • States: $s_{t}$ is defined as the dialogue history of previous $t$ turns, containing dialog act and agent/user utterances
    $h=\left(a_{0}^{\text{agent }},\mathrm{u}_{1}^{\text{agent }},\mathrm{u}_{1}^{\mathrm{usr}},\ldots,a_{t-1}^{\text{agent }},\mathrm{u}_{t}^{\text{agent }},\mathrm{u}_{t}^{\mathrm{usr}}\right)$ (3)
    , where $\mathrm{u}_{i}^{\text{agent }},\mathrm{u}_{i}^{\mathrm{usr}}$ are the utterances of the agent and user, respectively at the $i-$th turn.
  • Transitions: It "represents the dialogue state updates according to stochastic responses from the user to the agent." (Wang et al., 2020)
  • Rewards: it represents the immediate feedback of a desired conversational outcome, such as in-process, success, or failure of persuading a user to donate to a charity
2.6 Discussion
Accessibility of Action-State Transition
In environments under deterministic transitions (e.g., dialog, code generation), the next state $s^{\prime}$ can be directly derived based on the selected action. In a dynamic environment, $s^{\prime}$ can be either sampled over the probability distribution or generated from $\text{lmpr}_{\text{transition}}$. Section 4 will demonstrate how this property affects search procedures.
Overhead of Using MDP Definition
Although the comprehensive definition provides a unified interface to discuss LIS frameworks, it increases the overhead when applied to graph traversal tasks, since several defining characteristics of an MDP are not necessary, e.g., state transitions and explicit definitions of actions.
Why Is There No Previous Work Defining Reasoning Tasks as MDPs?
Typical MDPs are often defined for decision-making models which can only handle the tasks whose action space is constrained and finite. However, general-purpose models like LLMs naturally deal with infinite or/and hard-to-define action space, since LLMs can infer plausible actions with world knowledge and commonsense.
Do Tasks Enable Action Undoing and State Back-Up?
Some environments allow going back up to an earlier step after executing a sequence of actions (e.g., reasoning tasks), while other tasks may not (such as robotic tasks). This property is particularly important to discuss the applicability of LLM-integrated search methods. For environments under such property, the LIS agent can feel free to simulate future states for planning without worrying that the change of environments is irreversible. Details will be discussed in § 5).
3 LLM-Profiled Roles
Following standard reinforcement learning terminology (Sutton & Barto, 2018), an agent designed to solve Markov Decision Processes (MDPs) typically incorporates the following components:
  • Policy $\pi(a_{t}\mid g,s_{t})$: Determines the action $a_{t}$ to take given the current state $s_{t}$.
  • Value Function $V^{\pi}(s)\mapsto R$: Estimates the expected return of state $s$ under policy $\pi$.
  • Transition Model $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$: Represents the dynamics of the environment, predicting the next state $s_{t+1}$ given the current state $s_{t}$ and action $a_{t}$.
These definitions are broadly applicable across different agent designs. In this work, we adapt them to LLM-based search and focus on how to profile LLMs to work as/for these agentic components.
Background of LLM-Profiled Policy, Evaluator and Transition Model
This section outlines the implementation of the three core components using three types of LMPRs. These roles are defined by Li (2024) as the LLM-profiled policy ($\text{lmpr}_{\text{policy}}$), evaluator ($\text{lmpr}_{\text{eval}}$), and transition model ($\text{lmpr}_{\text{transition}}$). For brevity, these notations are commonly adopted throughout this work.
While prior studies such as Spiegel et al. (2024) and Feng et al. (2024) explored these LMPRs primarily in theoretical contexts and toy environments for reinforcement learning, this section extends these ideas by presenting detailed implementations in real-world tasks.
Presentation of Prompting Examples
To illustrate how LLMs are configured for different LMPR roles, we provide prompting examples throughout the paper. Model outputs are visually distinguished using shadow boxes for clarity. For example:
An output from LMPR
To maintain brevity, placeholders enclosed in angle brackets (e.g., <demos> for few-shot demonstrations and <task desc.> for task descriptions) are used to represent verbal components within prompts.
Outputs Example Works
$\text{lmpr}_{\text{naive\_policy}}$ $a_{t}$ Xie et al. (2023)
$\text{lmpr}_{\text{reasoning\_policy}}$ Reasoning, $a_{t}$ ReAct (Yao et al., 2023b)
$\text{lmpr}_{\text{batch\_policy}}$ $a_{t}^{1},a_{t}^{2},\ldots,a_{t}^{N}$ Yao et al. (2023a)
Table 3: LLM-Profiled Policy. Note that what we really require is only $a_{t}$. $N$: sample size; $T$: length of action sequence.
3.1 LLM-Profiled Policy (LMPP)
$\text{lmpr}_{\text{naive\_policy}}$
Given the observation $o_{t}$, $\text{lmpr}_{\text{naive\_policy}}$ directly generates the next action $a_{t}$.
$\text{lmpr}_{\text{reasoning\_policy}}$
To generate $a_{t}$, this policy first produces a complete reasoning path that explains or justifies the generation of $a_{t}$. The reasoning path serves as an explicit intermediate step, enhancing interpretability and illuminating the decision-making process for $a_{t}$.
$\text{lmpr}_{\text{react\_policy}}$
In contrast to $\text{lmpr}_{\text{reasoning\_policy}}$, $\text{lmpr}_{\text{react\_policy}}$ separates the reasoning step and the action-generation step into distinct inference passes. Each pass corresponds to an uninterrupted generation session. The reasoning text may include a planning path (e.g., $\tilde{a}_{t+1},\ldots,\tilde{a}_{T}$), but only $\tilde{a}_{t}$ is used for search. Another distinguishing feature is the more autonomous behavior of this policy, which does not strictly adhere to a fixed reasoning-then-acting sequence. Instead, it can dynamically alternate between reasoning and acting steps, such as reasoning-acting-acting. For example:
Your task is to: put a cool tomato in microwave.
think: To solve the task, I need to find a tomato, then cool it with the fridge, and finally put it in the microwave. <more thoughts>
OK.
> go to countertop 1
<observation>
> go to countertop 2
Prompting Example 1
The term “react” is attributed to the work of ReAct (Yao et al., 2023b). However, in their formulation, each thought is not explicitly treated as an action; instead, only tool invocations are considered actions in reasoning tasks. This distinction highlights the broader applicability of $\text{lmpr}_{\text{react\_policy}}$ in our definition.
$\text{lmpr}_{\text{batch\_policy}}$
This policy generates a batch of actions simultaneously, in contrast to other $\text{lmpr}_{\text{policy}}$ approaches, which generate actions step-by-step. $\text{lmpr}_{\text{batch\_policy}}$ can sample all actions in one pass. The sampling procedure is specified in Section 4. As noted by Yao et al. (2023a), this avoids duplication in constrained action spaces, such as selecting a word in a crossword puzzle. By leveraging a global view of the action space, $\text{lmpr}_{\text{batch\_policy}}$ improves efficiency and coherence in tasks.
3.2 LLM-Profiled Transition Model (LMPT)
$\text{lmpr}_{\text{transition}}$ predicts outcomes according to LLMs’ internal knowledge. The profiling can be categorized as generating: 1) Full state: The final goal is to return a full state/observation at the current step, as exemplified in Example 3.2.
<profile information>
[STATE 0] I have that, the white block is clear, the cyan block is clear, <more detail>
[ACTION] Pick up the brown block.
[CHANGE]
The hand was empty and is now holding the brown block, the brown block was on the table and is now in the hand, and the brown block is no longer clear. [STATE 1] I have that, the white block is clear, the cyan block is clear, <more detail>
Prompting Example 2
2) Partial observation: The partial observation would be further processed to form the full state. One obvious task is reasoning via QAs.
Given a question, please decompose it into sub-questions. For each sub-question, please answer it in a complete sentence, ending with "The answer is". When the original question is answerable, please start the subquestion with "Now we can answer the question:
<few shot demos>
Question 1: James writes a 3-page letter to 2 different friends twice a week. How many pages does he write a year?
Question 1.1: How many pages does he write every week?
Answer 1.1: James writes a 3-page letter to 2 different friends twice a week, so he writes 3 * 2 * 2 = 12 pages every week. The answer is 12.
Prompting Example 3
Table 4: LLM-profiled evaluators $\text{lmpr}_{\text{eval}}$. The columns of “$V$?” and “$Q$?” indicate whether the LMPR configuration can work as state value function and action value function, respectively.
Prompting Tasks Outputs V? Q? Example Works
$\text{lmpr}_{\text{eval1}}$ Binary/Multi-class
Classification
Discrete values mapped by $\text{lmpr}_{\text{eval}}$ generations RAP (Hao et al., 2023), ToT (Yao et al., 2023a), Koh et al. (2024), LATS (Zhou et al., 2024a)
$\text{lmpr}_{\text{eval2}}$ Binary/multi-class Classification Logits of $\text{lmpr}_{\text{eval}}$ generations RAP (Hao et al., 2023), Tree-BeamSearch (Xie et al., 2023)
$\text{lmpr}_{\text{eval3}}$ Multi-choice QA Choices of top-N actions ToT (Yao et al., 2023a)
$\text{lmpr}_{\text{eval4}}$ Classification (Implicit) Logits of given continuation RAP (Hao et al., 2023)
$\text{lmpr}_{\text{eval5}}$ Multi-choice QA Logits of given continuation (choices) Kadavath et al. (2022), Xie et al. (2023)
$\text{lmpr}_{\text{policy\&eval1}}$ NA Logits of $\text{lmpr}_{\text{policy}}$ generations RAP (Hao et al., 2023), Tree-BeamSearch (Xie et al., 2023)
$\text{lmpr}_{\text{policy\&eval2}}$ NA Self-consistency scores of $\text{lmpr}_{\text{policy}}$ generations LATS (Zhou et al., 2024a)
$\text{lmpr}_{\text{policy\&eval3}}$ NA Consistency of $\text{lmpr}_{\text{policy}}$-generated steps and candidate states LLM-A* (Meng et al., 2024)
$\text{lmpr}_{\text{policy\&eval4}}$ NA Q-Values based on a stroing $\text{lmpr}_{\text{policy}}$ Q* (Wang et al., 2024a)
$\text{lmpr}_{\text{transition\&eval1}}$ NA Logits of $\text{lmpr}_{\text{transition}}$ generations RAP (Hao et al., 2023)
$\text{lmpr}_{\text{transition\&eval2}}$ NA Logits of $\text{lmpr}_{\text{transition}}$ generations /
3.3 LLM-Profiled Evaluator (LMPE)
LLMs can serve as flexible evaluators (LMPEs) by leveraging their generative and probabilistic capabilities. We propose categorizing these evaluators along three key dimensions:
  • Task Formulation: Whether the evaluation is a binary classification, multi-choice QA, or a free-form judgment influences how the LLM’s output or logits can be interpreted. These tasks are always formulated in LLMs’ system-level prompts.
  • State vs. Action Evaluation: This is analogous to state/action-state value functions in reinforcement learning. Depending on whether the evaluator is assessing a static state $s_{t}$ or a transition $(s_{t},a_{t})$, the LLM must parse different context inputs to provide a valid judgment.
  • Output Types: Unlike $\text{lmpr}_{\text{policy}}$ or $\text{lmpr}_{\text{transition}}$, the output of an evaluator $\text{lmpr}_{\text{eval}}$ can produce not only text-based outputs (continuous text or discrete labels) but also raw logits and self-consistency scores. This flexibility allows for nuanced scoring and confidence measurement that can be mapped to discrete classes or continuous values.
  • Use of Reasoning: As LMPPs, some reasoning techniques can be generalized to LMPEs.
As summarized in Table 4, various works have shown that LLMs configured under these dimensions can support diverse evaluation goals. The paragraphs below illustrate the four dimensions through concrete prompts and output examples.
Task Formulations
Three types are commonly used for evaluation. 1) Binary/Multi-class Classification: As shown below, a prompt can explicitly request a binary judgment, e.g., yes/no or failure/success:
<prompt>
Status: “failure”
Prompting Example 4
In some cases, more fine-grained judgments are required with multiple labels. For example, the “status” tag is defined for LMPE to indicate partial successes in Koh et al. (2024):
<prompt>
Status: “failure”
On the right track to success: “yes”
Prompting Example 5
This can be considered as multi-class classification, where “yes” yields an intermediate class. A more direct multi-class classification is specified in the example below, where the prompt assesses whether a given set of numbers can reach 24:
Evaluate if given numbers can reach 24 (sure/likely/impossible)
10 14
10 + 14 = 24
sure
1 3 3
1 * 3 * 3 = 9
(1 + 3) * 3 = 12
1 3 3 are all too small
impossible
Prompting Example 6
2) Multi-choice QA: This is often advantageous when directly scoring an action/state is difficult to compute in contrast to comparing multiple candidates. For example, it is difficult to judge whether a give passage is coherent, while it is easy to judge whether Passage A is more or less coherent than Passage B. Another way is to implicitly compare different solutions via voting through self-consistency scores of $\text{lmpr}_{\text{policy}}$, which belongs to the next formulation type. 3) No Explicit Evaluator Definition: Evaluation can be inferred from the $\text{lmpr}_{\text{policy}}$’s generative process itself. In such cases, no separate system-level prompt is required for task formulation. Likewise, $\text{lmpr}_{\text{transition}}$ can be used for evaluation. This LLM-based evaluation will be detailed from the perspective of output types.
State vs. State-Action Function
1) State-Value Evaluator: A state-based evaluator accepts $s_{t}$ as its input to produce a judgment:
discrete $\text{discrete judge}=\text{lmpr}_{\text{eval}}(s_{t})$ (4)
Example 3.3 is one of the example. 2) State-Action Evaluator: the evaluator assesses whether taking action $a_{t}$ is appropriate at the current state $s_{t}$:
discrete $\text{discrete judge}=\text{lmpr}_{\text{eval}}(s_{t},a_{t})$ (5)
This setup is exemplified in Example 3.4, where the new sub-question ($a_{t}$)’s usefulness depends on the prior state ($s_{t}$). Another example of BlocksWorld from Hao et al. (2023):
[STATE]
As initial conditions I have that, the blue block is clear, the orange block is in the hand, the red block is clear, the hand is holding the orange block, the red block is on top of the yellow block, the blue block is on the table, and the yellow block is on the table. My goal is to have have that the red block is on top of the yellow block and the orange block is on top of the blue block.
[ACTION]
stack the orange block on top of the red block
[EVALUATION]
bad
Prompting Example 7
Outputs
Finally, the outputs of $\text{lmpr}_{\text{eval}}$ can take one of several forms, depending on how we wish to interpret or utilize the evaluator’s opinion:
  1. 1.
    Mapping $\text{lmpr}_{\text{eval}}$ generation to discrete values: For instance, “impossible” or “Yes” may be mapped to numeric scores (0 or 1) in Quotes 3.3 and 3.4.
  2. 2.
    Using logits of $\text{lmpr}_{\text{eval}}$: The probability of generating a specific token (e.g., “impossible”) can serve as the confidence score.
  3. 3.
    Using logits of given continuations: Rather than having the LLM generate the evaluation tokens, one can provide the exact sequence to be evaluated (e.g., “good” in Quote 3.3 or “(A) good”). The log probability of each token is then summed to indicate how well (i.e., how confidently) the model “accepts” that evaluation in the given context. A higher cumulative log-likelihood suggests that the LLM finds the provided evaluation more plausible. Moreover, multiple predefined options (e.g., “(A) Correct” vs. “(B) Incorrect”) can be separately fed in as continuations. The log probabilities of each option can then be compared as self-evaluation scores (Xie et al., 2023; Kadavath et al., 2022).
  4. 4.
    Using logits of $\text{lmpr}_{\text{policy}}$ or $\text{lmpr}_{\text{transition}}$: Alternatively, the evaluation can be derived from the policy or transition model’s token probabilities. This method can avoid additional inference steps by reusing existing logits.
  5. 5.
    Using self-consistency scores: By sampling multiple trajectories from $\text{lmpr}_{\text{policy}}$ or $\text{lmpr}_{\text{transition}}$, one can gauge confidence via how often a particular outcome (state or action) appears. More frequent outcomes can be assumed more likely (or better).
  6. 6.
    Comparing consistency of $\text{lmpr}_{\text{policy}}$-generated steps and candidate states: The actions/plan generated by actions can be compared to the candidate for evaluation. This is suitable for tasks with limited successor states.
  7. 7.
    Comparing Consistency of $\text{lmpr}_{\text{policy}}$-Generated Steps and Candidate States: The actions or plans produced by the policy can be compared against candidate states to assess consistency. This approach is particularly suitable for tasks with a limited number of successor states, e.g., sovling a maze (Meng et al., 2024).
  8. 8.
    Using a Stronger $\text{lmpr}_{\text{policy}}$ as a Proxy Optimal Policy to Approximate Q-Values: When rewards are only obtained at the terminal state, the Q-value can be approximated by discounting the rewards along the path. This approach assumes that all subsequent actions are optimal, which necessitates the use of a more robust $\text{lmpr}_{\text{policy}}$ as a proxy for the optimal policy.
Use of Reasoning
Similar to $\text{lmpr}_{\text{reasoning\_policy}}$, a reasoning process can be required before generating the final judgment. This reasoning process provides a logical justification to augment the evaluation. For example:
<prompt>
Thoughts: <your thoughts and reasoning process> Status: “failure”
Prompting Example 8
3.4 Discussion
Inference Cost of $\text{lmpr}_{\text{policy}}$
In practice, the overall computational cost follows the pattern
$\text{lmpr}_{\text{react\_policy}}\;>\;\text{lmpr}_{\text{reasoning\_policy}}\;>\;\text{lmpr}_{\text{naive\_policy}}.$
The gap between $\text{lmpr}_{\text{reasoning\_policy}}$ and $\text{lmpr}_{\text{naive\_policy}}$ arises from the additional output tokens produced for reasoning. More importantly, when commercial API is used, $\text{lmpr}_{\text{react\_policy}}$ exhibits an even higher cost because each separate reasoning or action-generation pass is effectively stateless with respect to the cached K–V pairs from previous passes, thereby preventing token-reuse optimizations.
Applicability of $\text{lmpr}_{\text{react\_policy}}$.
A central requirement for ReAct-style prompting ($\text{lmpr}_{\text{react\_policy}}$) is the availability of step-wise observations after each action. This imposes two prevalent scenarios:
  1. 1.
    Tasks relying on simulators: When direct interaction with the real environment is impractical (e.g., actions on tasks are irreversible), a simulator can be substituted to generate the observation following each action. For instance, an LLM-based simulator ($\text{lmpr}_{\text{transition}}$) might use commonsense knowledge to model environmental responses (e.g., turning on a water tap in a sealed sink). However, such simulators are unsuitable for tasks involving external or private data—like querying proprietary databases or retrieving up-to-date information—since an LLM’s internal knowledge typically cannot replicate these data sources.
  2. 2.
    Action-reversible tasks. Certain problems can be retried or backtracked, allowing the agent to iteratively act, observe, and refine its actions, as discussed in Section 2. In Section 5, for example, LLM-based search frameworks such as LATS (Zhou et al., 2024a) leverage this property when interacting with real environments across multiple search steps to perform monte-carlo simulation.
Risk of $\text{lmpr}_{\text{value}}$
Although $\text{lmpr}_{\text{value}}$ can effectively evaluate state or action quality, two challenges stand out:
  1. 1.
    Mediocre discrimination abilities: As shown by Chen et al. (2024b), using logits as dense rewards (e.g., in $\text{lmpr}_{\text{policy\&eval}}$ or $\text{lmpr}_{\text{transition\&eval}}$) can reveal that many open-source LLMs struggle to reliably distinguish “good” from “bad” examples. [^1] GPT-4 turbo was the most advanced model at the time of evaluation.
  2. 2.
    In-Context Reward Hacking (ICRH): According to Pan et al. (2024), an LLM evaluator ($\text{lmpr}_{\text{value}}$) may attempt to “explain away” negative feedback by globally altering its reasoning and actions, potentially violating constraints. For example, to fix an InsufficientBalanceError, the LLM might suggest unauthorized money transfers from other accounts, thus compromising safety or policy compliance.
Not All Generation with “Reasoning” is Truly Augmented.
By design, LLMs generate tokens in an auto-regressive manner, meaning earlier tokens are not influenced by later ungenerated tokens. Hence, although reasoning tokens after actions (or evaluation) can make the model outputs more interpretable, they do not always alter subsequent decisions or evaluations. In Xie et al. (2023), for instance, a chain of thoughts
$a_{t},\;\tilde{a}_{t+1},\;\ldots,\;\tilde{a}_{T}$
is produced, where $\tilde{a}_{t+1},\ldots,\tilde{a}_{T}$ are “unrecorded” actions. Crucially, $a_{t}$ is unaffected by any future $\tilde{a}$ tokens, making this effectively a naive policy rather than a true reasoning-augmented approach.
Similarly, consider the evaluator in Example 3.4:
Given a question and some sub-questions, determine whether the last sub-question is useful to answer the question. Output ’Yes’ or ’No’, and a reason.
Question 1: Four years ago, Kody was only half as old as Mohamed. If Mohamed is currently twice as 30 years old, how old is Kody?
Question 1.1: How old is Mohamed?
Question 1.2: How old was Mohamed four years ago?
New question 1.3: How old was Kody four years ago?
Is the new question useful?
Yes. We need the answer to calculate how old is Kody now.
Prompting Example 9
Although the model’s output includes a short “reason,” that intermediate reasoning does not necessarily inform the generation of ‘Yes’ or ‘No’.
Table 5: Examples of combining LLM-Profiled Evaluators with heuristics. $\|s=s_{g}\|_{1}$ indicate whether the agent reaches the goal state at $s$.
Task Value
Reasoning-QA in RAP (Hao et al., 2023) $\text{lmpr}_{\text{eval1}}$ + $\text{lmpr}_{\text{policy\&eval1}}$
Game (Blockworld) in RAP (Hao et al., 2023) $\text{lmpr}_{\text{eval4}}+\text{lmpr}_{\text{eval2}}+\|s=s_{g}\|_{1}$ (weights ignored)
Graph traversal in LLM-A* (Meng et al., 2024) Euclidean distance to the next state + $\text{glm}_{\text{eval}}$
Reasoning in Xie et al. (2023) $\text{lmpr}_{\text{value5}}+\text{lmpr}_{\text{policy\&eval1}}$
Reasoning-QA/Code Gen/Web Nav. in LATS (Zhou et al., 2024a) $\text{lmpr}_{\text{value1}}+\text{lmpr}_{\text{policy\&eval2}}$
4 Search Procedures
This section presents the reusable search procedures applied across various frameworks, including both non-LMPR-specific and LMPR-based procedures. Unlike Section 3, which focused on configuring LMPRs, here we demonstrate how these LMPRs are integrated into the operational processes. However, some content may overlap slightly for coherence. Table 6 summarizes the dependencies between these search procedures and the LMPR components.
Table 6: Overview of dependencies in search procedures. Sampl.: Sampling; Exp.: Expansion; Eval.: Evaluation; Sel.: Selection; Sim.: Simulation; Backprop.: Backpropagation.
(a) Dependency of first-order procedures on LMPRs.
LMPP LMPE LMPT
LMPP Sampl.
LMPE+ Eval.
LMPT Sim.
Value-Based Sel. maybe
Single-Step UCT Sel.
Exhaustive Action Retrieval
(b) Dependency of higher-order procedures on LMPR-based, first-order procedures.
LMPP Sampl. LMPE+ Eval. LMPT Sim.
Value-Based Sel. maybe
LMPP Exp. maybe
Path Sim. maybe maybe
MCTS Sel.
MCTS Backprop.
Search Nodes: Integrating States, Action, and Rewards
In this section, we shift our focus to search and clarify how the fundamental search “node” is defined with respect to states and actions. Some methods (e.g., ToT (Yao et al., 2023a)) treat a node as a particular state in a search tree, with transitions determined by the actions taken. To ensure generality, we unify states, actions, and even their estimated values (or rewards) in a single node structure (e.g., RAP (Hao et al., 2023)), facilitating partial expansions or multi-step lookahead. To align with object-oriented design, we represent a node as $n$ with attributes $n.\text{action}$, $n.\text{state}$, $n.\text{parent}$, and $n.\text{val}$ , representing the action, state, parent node, and value, respectively.
4.1 First-Order Procedures
First-order procedures operate independently, without relying on other procedures. They serve as the foundational components upon which more complex procedures are built, ensuring a modular and scalable framework for LLM-based search operations. The first three are based on LLM-Profiled Policy (LMPP), evaluator (LMPE), and transition model (LMPT), respectively, while others not necessarily depend on LMPRs.
LMPP Sampling
The sampling procedure involves generating multiple actions (assuming $N$ actions) for a given state $s_{t}$. Generally, there are two approaches to sampling actions: one based on generating actions sequentially using the single-action policy ($\text{lmpr}_{\text{policy}}$), and another based on generating all actions simultaneously using the batch policy ($\text{lmpr}_{\text{batch\_policy}}$). These approaches are detailed in Procedures 1 and 2, respectively.
List of procedures 1 LMPP Sampling: Sample Actions One at a Time
1:procedure Sample_LMPP_One_At_A_Pass($s_{t},N$)
2:    $\mathbf{a}_{t}\leftarrow\{\}$
3:    for $i\in\{1,\ldots,N\}$ do
4:         $a_{i}\sim\text{lmpr}_{\text{policy}}\left(a\mid s_{t}\right)$
5:         $\mathbf{a}_{t}\leftarrow\mathbf{a}_{t}\cup\{a_{i}\}$
6:    end for
7:    return $\mathbf{a}_{t}$
8:end procedure
List of procedures 2 LMPP Sampling: Sample All Actions at Once
1:procedure Sample_LMPP_All_At_Once($s_{t},N$)
2:    $\mathbf{a}_{t}\sim\text{lmpr}_{\text{batch\_policy}}\left(a\mid s_{t},N\right)$
3:    return $\mathbf{a}_{t}$
4:end procedure
LMPE+ Evaluation
LMPEs can be used to estimate the value (or reward) of a state. The evaluator’s output—whether in textual or numerical form—can then be combined with rule-based heuristics to refine the overall assessment. For instance, Table 5 illustrates how numeric outputs from LMPEs are incorporated into a heuristic that balances both LLM-based scoring and domain-specific constraints. Such integrated approaches are particularly relevant when neither pure heuristic nor pure LLM-based evaluation alone is sufficient for robust decision-making. Based on the input, the procedure can be defined as V_Eval for taking a state as input and Q_Eval for taking both a state and an action as input.
LMPT Simulation
This procedure is straightforward: given the current state $s_{t}$ and an action $a_{t}$, the LMPT directly outputs the next state $s_{t+1}$. Unlike LMPP sampling (which may loop through multiple actions) or LMPE+ Evaluation (which may incorporate additional heuristics), no further processing or components are involved.
Single-Step UCT Selection
The objective of UCT selection is to choose an action that balances exploitation and exploration. This is captured by the Upper Confidence Tree (UCT) formula:
$\operatorname{UCT}\left(s,a\right)=Q(s,a)+w\sqrt{\frac{\ln N(s)}{N(c(s,a))}},$ (6)
where $c$ is a constant controlling exploration, $N(s)$ is the number of times state $s$ has been visited, and $N(s,a)$ is the number of times action $a$ has been selected under $s$. A related variant, Predictor Upper Confidence Tree (PUCT), incorporates the prior probability $P(s,a)$ into the exploration term to further guide the action selection.
$\operatorname{PUCT}\left(s,a\right)=Q\left(s,a\right)+\mathrm{c}P(s,a)\frac{\sqrt{N\left(s\right)}}{1+N\left(s,a\right)}$ (7)
$P(s,a)$ is normally a domain-specific predictor. Based on the two estimates, the procedure is just to iterate over the given actions $A$, along with their values $Q$, and extract the one with the highest value, as summarized in Procedure 3.
List of procedures 3 Single-Step UCT Selection
1:procedure UCT_Select($s,A$)
2:    $a^{*}=\arg\max_{a\in A}\left[\operatorname{UCT}(s,a)\right]$
3:end procedure
4:procedure PUCT_Select($s,A$)
5:    $a^{*}=\arg\max_{a\in A}\left[\operatorname{PUCT}(s,a)\right]$
6:end procedure
Exhaustive Action Retrieval
When the action space is small and well-defined (e.g., in BlockWorld), all possible actions can be retrieved exhaustively. This procedure is primarily used to facilitate the subsequent expansion or simulation steps.
4.2 Higher-Order Procedures
Value-Based Selection
The first type is top-k selection. The top $k$ states or actions are picked from a large pool of candidates based on their estimated values. A state-value function $V(s^{\prime})$ or an action-value function $Q(s,a)$ is used to generate values. Commonly, they are implemented by LMPE+ evaluation. The detail is illustrated in Procedures 4. Note that the if statement for value assignment also allows specialized ways to assign values without necessarily relying on $n^{\prime}.\text{state}$.
List of procedures 4 TopK-Based Selection
1:procedure TopK_Select($N,k$)
2:    for each node $n^{\prime}\in N^{\prime}$ do
3:         if $n^{\prime}.\text{val}$ is uninitialized then
4:             $n^{\prime}.\text{val}\leftarrow\text{ValueFunc}(n^{\prime}.\text{state})$
5:         end if
6:    end for
7:    $N^{*}\leftarrow\arg\operatorname{top}_{k}\bigl{\{}n^{\prime}.\text{val}\mid n^{\prime}\in N^{\prime}\bigr{\}}$
8:    return $N^{*}$
9:end procedure
Once $k=1$, the $\arg\operatorname{top}_{k}\bigl{\{}n^{\prime}.\text{val}\mid n^{\prime}\in N^{\prime}\bigr{\}}$ reduces to $\arg\max_{n^{\prime}\in N^{\prime}}n^{\prime}.\text{val}.$
Another type is threshold-based selection. Here, the procedure repeatedly samples one action from the current node’s state, simulates the new state, and evaluates its value. If the value surpasses a threshold $\theta$, the procedure returns the newly created node; otherwise, it continues sampling.
List of procedures 5 Threshold-Based Selection
1:procedure ThresholdSelect($n,\theta$)
2:    while true do
3:         $a\leftarrow\text{Sample\_Action}(n.\text{state})$
4:         $s^{\prime}\leftarrow\text{Simulate}(n.\text{state},a)$
5:         $v\leftarrow\text{ValueFunc}(s^{\prime})$
6:         if $v\geq\theta$ then $\triangleright$ Value exceeds threshold; accept this node
7:             Instantiate a new node $n^{\prime}$
8:             $n^{\prime}.\text{state}\leftarrow s^{\prime}$
9:             $n^{\prime}.\text{action}\leftarrow a$
10:             $n^{\prime}.\text{val}\leftarrow v$
11:             return $n^{\prime}$
12:         end $\triangleright$ Otherwise, continue sampling another action
13:    end while
14:end procedure
LMPP Expansion
This expansion procedure adds one or more child nodes under the given leaf node(s), typically visualized at the next depth level in a tree search. The procedure is based on the following components:
  1. 1.
    LMPP Sampling: This step provides the $n.\text{action}$ attribute for each node, while leaving $n.\text{state}$ and $n.\text{val}$ uninitialized.
  2. 2.
    Simulating States Based on Sampled Actions: To enable further expansion, the node’s state ($n.\text{state}$) must be generated. This can be accomplished via simulators (e.g., using LMPT simulation), direct action execution, or even hard-coded methods (e.g., simply concatenating actions).
  3. 3.
    (Optional) Expanding Multiple Nodes Simultaneously: In many search strategies (such as beam search with a beam size greater than 1 or breadth-first search), multiple leaf nodes are expanded concurrently.
A general form of the expansion procedure is specified in Procedure 6.
List of procedures 6 Expansion
1:procedure Expand_LMPP($N_{t}$)
2:    $N_{t+1}\leftarrow\{\}$
3:    for each node $n_{t}\in N_{t}$ do
4:         $A_{t}\leftarrow\text{{Sample\_LMPP}}(n_{t}.\text{state},k)$
5:         for each action $a_{t}\in A_{t}$ do
6:             Instantiate a new node $n_{t+1}$ with $n_{t+1}.\text{action}\leftarrow a_{t}$
7:             $n_{t+1}.\text{state}\leftarrow\text{{Simulate}}(n_{t}.\text{state},a_{t})$
8:             $N_{t+1}\leftarrow N_{t+1}\cup n$
9:         end for
10:    end for
11:    return $N_{t+1}$
12:end procedure
Here, Sample_LMPP refers to Procedure 1 or Procedure 2. The total number of $N_{t+1}$ equals the product of the number of nodes in $N_{t}$ and the number of actions $k$ sampled per node.
Path Simulation
Procedure 7 demonstrates the overall process for rolling out a path until either a goal state or a terminal state is reached. In this procedure, the following steps are executed sequentially and iteratively:
  1. 1.
    Sampling: At each step, actions are sampled from the current state using either LMPP sampling or a random sampling method.
  2. 2.
    Value-Based Selection: TopK_Select is applied with only the highest-valued node preserved (i.e., $k=1$), since only a single path is rolled out. The evaluation can be performed using LMPE+ evaluation or a heuristic-only approach.
    3) Single-Step Simulation: The next state is generated either by either using LMPT simulation, employing domain-specific simulators, or directly executing the selected action in the environment to obtain the subsequent state (or observation).
List of procedures 7 Path Simulation
1:procedure Simulate($n_{t}$)
[…]
3:    while $n_{t}.\text{state}$ is not terminal do
4:         $A\leftarrow\text{{Sample\_Action}}(n_{t}.\text{state})$
5:         $a^{*}\leftarrow\text{{Q\_Select}}(n_{t}.\text{state},A^{\prime},k=1)$
6:         $s\leftarrow$ Simulate($s,a^{*}$)
7:         path $\leftarrow\text{path}\cup s$
8:    end while
9:    return path
10:end procedure
MCTS Selection
However, during MCTS for planning, before going to the expansion phase, Procedure 3 (One-step UCT Selection) should be used multiple times to traverse from $s_{0}$ to a leaf node $s_{\text{leaf}}$.
MCTS Backpropagation - Value Update
After each simulation returns a reward $r$, update the Q value as:
$Q_{\text{new }}=\frac{r+Q_{\text{old }}\cdot\text{Count}_{\text{new }}}{\text{Count}_{\text{new }}},$ (8)
  • $r$, depending on the task, can be a reward at the terminal state. In some cases,it can be an aggregated one, if each simulation step yields a reward. Specifically, if rewards $r_{t}$ are discounted by $\gamma$, then the final sample reward $r$ for backpropagation is:
    $r=G=\sum_{t=0}^{T-1}\gamma^{t}r_{t}.$
    In many implementations (e.g., RAP (Hao et al., 2023)), the step-wise rewards are obtained via LMPE+ or a heuristic evaluation during path simulation. These rewards from simulated nodes are then employed to update the Q values for non-simulated nodes, including the leaf nodes and those above.
  • $Q_{\text{old}}$ are the previous $Q$-value.
  • $\text{Count}_{\text{new}}$ is the total visit count after the current update.
During some implementations (Hao et al., 2023), each reward $r\in R_{\text{cum}}$ propagating from the terminal node can be stored in a list $R_{\text{cum}}$ , and average them when used. The procedure of value estimate, as summarized in Table 5, provides the initialized values for new actions or resulting states.
MCTS Backpropagation - Visit Update
Except for the estimated value $Q(s,a)$, the backpropagation also updates the visit count for every state on the path from the root to the leaf and each edge $(s,a)$ along that path, denoted as $\text{Count}\left(s_{t}\right)$ and $\text{Count}\left(s_{t},a_{t+1}\right)$, respectively. The increase in $\text{Count}(s,a)$ (the action-level visit count) reduces the exploration bonus for $a$ in future selections, thus making ( $\mathbf{s},\mathbf{a}$ ) slightly less likely to be chosen purely for exploration next time, assuming the same or lower estimated value. Assuming $\text{Count}\left(s_{t},a_{t+1}\right)->\text{Count}\left(s_{t+1}\right)$ in some deterministic environments, only $\text{Count}\left(s_{t+1}\right)$ is tracked. For the same reason, $Q$ values can be attached as an attribute of the state node. An example is the implementation of RAP (Hao et al., 2023)  [^2] https://github.com/maitrix-org/llm-reasoners/blob/main/reasoners/algorithm/mcts.py.
11footnotetext: multimodal22footnotetext: fine-tuning
Table 7: Search-based frameworks for LLM inference. EAR means Exhaustive Action Retrieval. We use “*” to indicate workshop publication. Many MCTS-specific procedures are excluded, except for UCT Selection (UCT Sel.) to distinguish PUCT and UCT. Simulation differentiates LMPT from environment (Env.) simulation.
Resemble LMPP Exp. LMPE+ Eval. LMPT Sim. EAR UCT Sel. Published Date
PG-TD (Zhang et al., 2023) MCTS LMPT PUCT Mar 2023 (ICLR2023)
ToT (Yao et al., 2023a) BFS (B for Breath); DFS May 2023 (NIPS2023)
Xie et al. (2023) Beam Search May 2023 (NIPS2023)
RAP (Hao et al., 2023) MCTS LMPT UCT May 2023 (EMNLP2023)
GDP-ZERO (Yu et al., 2023) MCTS LMPT Oct 2023 (EMNLP2023)
LATS (Zhou et al., 2024a) MCTS Env. UCT Oct 2023 (ICML2024)
LLM-MCTS (Zhao et al., 2023) MCTS Simulator PUCT May 2023 (NIPS2023)
LLM-A* (Meng et al., 2024) A* Jun 2024 (EMNLP2024)
Q* (Wang et al., 2024a) A* optional Jun 2024
Koh et al. (2024) BFS (B for Best) 1 Env. Jul 2024
PathFinder (Golovneva et al., 2023) Beam Search Dec 2023 (NeurIPS2023*)
5 Frameworks Based on Search Algorithms
This section summarizes how different frameworks utilize search algorithms, leveraging the LMPRs and search procedures introduced in Table 7. Note that, some MCTS-specific procedures (e.g., MCTS selection, and MCTS backpropagation) are not elaborated in the table. UCT selection is highlighted to distinguish between PUCT and UCT variants; simulation is included to clarify whether LMPT or environment-based simulation or simulator is used. Below, we discuss perspectives that are not fully captured in Table 7.
5.1 Beam Search
Beam search can be employed for reasoning tasks through two procedures iteratively:
  1. 1.
    LMPP Expansion: Each set of beam nodes $N_{t}$ is passed to the LMPP expansion procedure. Internally, Sample_LMPP calls either Sample_Actions_One_At_A_Pass or Sample_Actions_Batch, while Simulate can be a simple concatenation transition: each node $n_{t}\in N_{t}$ has its parent state $n_{t}.\text{parent}$ expressed as a sequence of actions $(a_{1},\dots,a_{t-1})$, an action $a_{t}$ assigned to $n_{t}.\text{action}$, and the new node’s state ($n_{t}.\text{state}$) as $(a_{1},\dots,a_{t-1},a_{t})$. The resulting set of expanded nodes is denoted $N_{t+1}^{\text{sample}}$.
  2. 2.
    TopK-Based Selection: From $N_{t+1}^{\text{sample}}$, a value-based selection procedure (e.g., Procedure 4) is applied to pick the top-$k$ nodes (the beam size). This subset is returned as $N_{t+1}$.
  3. 3.
    The process repeats until reaching a terminal state.
Below are two frameworks that illustrate how beam search is adapted:
Xie et al. (2023)
  • Uses a value function implemented by $\text{lmpr}_{\text{value5}}+\text{lmpr}_{\text{policy\&eval1}}$.
PathFinder (Golovneva et al., 2023)
  • Computes a summed similarity score as the value for each candidate node, comparing its state with those of other beams $N_{t+1}^{\text{sample}}$. The similarity function can be as simple as n-gram overlap.
5.2 Breadth-First Search
Tree-of-Thoughts (ToT) (Yao et al., 2023a)
  • Similar to beam search, breadth-first search (BFS) is performed by iteratively applying LMPP Expansion and TopK-Based Selection.
  • A key difference is that, in BFS, all nodes at depth $t$ undergo the same number of actions before expanding further levels. This enforces uniform depths across expansions.
5.3 Depth-First Search
Tree-of-Thoughts (ToT) (Yao et al., 2023a)
Yao et al. (2023a) also apply depth-first search (DFS) for LLM inference, relying on the LMPP Expansion and Threshold-Based Selection. Key points include:
  • Threshold-Based Selection: One action (node) is sampled at a time, but it is not compared with other nodes. Instead, once its value is evaluated by whether it surpasses a threshold, that path is followed to its conclusion.
  • LMPE Evaluation for Deadend: The system uses a deadend judgment to halt exploration of unpromising paths, which can be considered as another LMPE evaluation.
  • Backtracking: Upon reaching a deadend, the system reverts to an earlier node and continues exploring other previously expanded but untried branches.
  • Path Maintenance: Because of backtracking, the framework must track partial paths, whereas BFS or beam search only needs to maintain the selected nodes at each depth.
5.4 Best-First Search
Best-first search typically uses a heuristic function $h(s)$ to estimate how promising a state will reach the goal.
Koh et al. (2024)
  • Uses LMPP Expansion for the selected node or the initial root node, where actions are executed in the environment (web interface) to simulate the next states. The generated nodes are saved in $N$.
  • Employs TopK_Select (k=1) through LMPE+ evaluation on each node in saved nodes $n\in N$. The node value $n.\text{val}$ is derived from evaluating its parent’s state.
  • Continues until either the search tree reaches a specified budget $\beta$ or the state value exceeds a threshold $\theta$.
5.5 A*
A* is similar to best-first search but augments the heuristic $h(s)$ with the accumulated cost/utility $g(s)$ to reach a node from the start. The evaluation function is
$f(s_{t})=g(s_{t})+\lambda\,h(s_{t}),$ (9)
where $\lambda$ balances the two terms. Table 8 summarizes how two frameworks implement this formula differently:
$\operatorname{Agg}$ for $g(s)$ $R(s)$/$\operatorname{Cost}(s)$ $h(s)$ LMPE for $h(s)$
LLM-A*
(Meng et al., 2024)
$\sum$ $\operatorname{dist}(s_{0},s_{n})$ $\|s_{n}-s_{g}\|$ + $\|s_{n}-s_{\text{llm}}\|$ $\text{lmpr}_{\text{policy\&eval3}}$
Q*
(Wang et al., 2024a)
$\min$ , $\max$ , $\sum$, $\operatorname{Last}$ Human feedback, ground-truth, rules, LLM logits $\max_{a_{t}\in\mathcal{A}}Q^{*}\left(s_{t},a_{t}\right)$ $\text{lmpr}_{\text{policy\&eval4}}$
[…]
LLM-A* (Meng et al., 2024)
  • Designed for path-finding tasks (e.g., mazes).
  • $\bm{g(s_{t})}$: Computed incrementally as the path cost from $s_{0}$ to $s_{t}$. Formally,
    $g(s_{t})=\sum_{i=1}^{t}\operatorname{Cost}(s_{i}).$ (10)
  • $\bm{h(s)}$ under LMPE+ evaluation: The main modification beyond the typical A* is to integrate a LMPE to the $h(s)$. Specifically, $\text{lmpr}_{\text{policy\&eval3}}$ (see Table 4) is applied to evaluate the Euclidean distance from $s_{n}$ back to the recently visited $s_{\text{llm}}\in\text{lmpr}_{\text{policy}}$, along with the typical Euclidean distance between $s_{n}$ (expanded neighbour node) and $s_{g}$.
Q* (Wang et al., 2024a)
  • Targets reasoning and code-generation tasks.
  • $\bm{g(s_{t})}$: Aggregates rewards via
    $g(s_{t})=\operatorname{Agg}\left(\mathcal{R}(s_{1}),\ldots,\mathcal{R}(s_{t})\right),$ (11)
    where $\operatorname{Agg}\in\{\min,\max,\sum,\operatorname{Last}\}$. Under this definition, $\mathcal{R}\left(s\right)$ in $g(s_{t})$ can be calculated as human feedback, ground-truth, rules or via LMPE+ evaluation, depending on tasks.
  • $\bm{h(s)}$ optionally under LMPE+ evaluation: It is initialized as the optimal Q-value of $s_{t}$ over all the possible actions:
    $\max_{a_{t}\in\mathcal{A}}Q^{*}\left(s_{t},a_{t}\right)$ (12)
    Optionally, $\text{lmpr}_{\text{policy\&eval4}}$ introduces a stronger LMPP to approximate an optimal policy.
5.6 Monte Carlo Tree Search
Monte Carlo Tree Search typically involves selection, expansion, path simulation, and backpropagation. Most reviewed frameworks follow these steps. Below are some differences in notable frameworks:
RAP (Hao et al., 2023)
  • Applicable to BlocksWorld, Crosswords, and other reasoning tasks.
  • Sample_Action depends on whether the action space is finite and predefined (e.g., BlocksWorld, where exhaustive retrieval is used) or open-ended (e.g., Crosswords, where LMPP sampling is used).
LATS (Zhou et al., 2024a)
  • Targets tasks with reversible actions (e.g., certain reasoning problems).
  • Executes actions in the actual environment during path simulation, requiring actions to be reversible to allow repeated trials.
LLM-MCTS (Zhao et al., 2023)
  • Designed for robotic tasks.
  • Uses random sampling and a domain-specific simulator for path simulation, producing next states and rewards.
  • Adopts a domain-specific $P(s,a)$ in PUCT, which is derived from LMPP sampling to form an action distribution.
PG-TD (Zhang et al., 2023)
  • Specializes in code generation.
  • Treats the prior distribution $P(a\mid s)$ in PUCT as the LMPP token probabilities for the next token, given the partial program.
  • Uses LMPT simulation to generate partial programs, but internally adopts beam search to complete the path until a leaf node is reached. They refer to the whole process as “evaluation”).
GDP-ZERO (Yu et al., 2023)
  • Designed for goal-oriented dialogue.
  • Skips path simulation (the dialogue may terminate at any step), so LMPE evaluation is performed directly to produce a reward for MCTS backpropagation after expansion.
  • Still requires LMPT simulation for expansion. Particularly, two inference calls are needed to simulate both the system and user responses.
6 Discussion
In this section, we analyze how search frameworks for LLM inference deviate from traditional search algorithms, where and how they apply, and the resulting impact on performance and efficiency.
6.1 Deviations from Typical Search Algorithms
Beyond Finite, Fixed Search Space
Typical search algorithms, e.g., BFS (breath), deals with fixed search space and needs to keep track of all possibilities at each depth level, the large or infinite search space can lead to excessive memory consumption. LMPP sampling based on LLM priors makes BFS overcome this limitation.
Beyond Finite, Fixed Search Spaces
Classical BFS or DFS typically requires enumerating all successors at each depth, which can lead to massive memory usage in large or infinite search spaces. By contrast, LMPP sampling (based on LLM priors) can manage successor expansions more selectively, reducing the need to store every possibility at each level. Also, it is possible to handle tasks with an open and infinite action space.
Making “Uninformed” Search Informed
Traditionally, BFS and DFS are considered uninformed, exploring the search space without heuristics. LLM-based frameworks labeled as BFS or DFS often incorporate LMPP sampling or LMPE+ evaluation, effectively introducing heuristic knowledge from the LLM. Moreover, anticipating dead ends in DFS is feasible with LLM-based heuristics. Classic DFS only identifies dead ends when it exhausts neighbor nodes. With an LLM, the search can backtrack early if the model predicts an unpromising or “dead-end” scenario.
Compromised Optimality in A*
A* requires an admissible heuristic $h(s)$ to guarantee optimality. However, if $h(s)$ partially depends on a policy-generated state $s_{\text{llm}}$ from $\text{lmpr}_{\text{policy}}$, the heuristic may be overestimated. This breaks admissibility assumptions, meaning the final solution may no longer be strictly optimal.
Terminology Deviation for Heuristic Search: Cost vs. Return/Value
Many early applications of search algorithms, such as Best-First Search and A*, focused on minimizing quantities like distance, travel time, or energy expenditure. Generally, the cost-based evaluation can be considered as minimization objectives or eliminating “bad” states (or unpromising actions). However, in modern applications involving LLM-integrated search, such as web navigation or document retrieval, heuristics often reflect value estimates (positive polarity), especially when LLMEs tend to be defined to reflect the relevance or utility of states, which are better suited for maximization objectives or maintaining “good” states (or promising actions). We highlight this point for rigidity. However, these terms can be abstractly defined without indicating real-world semantics. For example, in game design, moving to a node which end up losing a life point can be given a negative reward, while a reward from gaining a key can be granted as a negative cost.
6.2 Applicability
Extending Uninformed Search to Dynamic Decision-Making
BFS and DFS were originally designed to explore predefined or easily generated state spaces (e.g., enumerating all children in a graph). This can be: 1) No Need for Transition: Traditional implementations of DFS and BFS do not require an explicit, computed transition model because they inherently rely on the graph structure where edges already define state transitions. For example, one task is to solve a maze. 2) Only Successor Function: In some applications, especially in implicit state-space search, a minimal form of transition function (or successor function) is still required to generate successors when the full graph is not explicitly available.
However, when applied to LLM inference, they can be adapted for: 3) Tasks That Require Dynamic Decision-Making: These tasks are based on state-dependent actions (as seen in planning or reinforcement learning), e.g., solving the Game of 24 and completing crossword puzzles.
Open-Loop vs. Closed-Loop Frameworks
This property is important to discuss the applicability of frameworks for planning. 1) Open-Loop (Offline): After executing $a_{t}$, the open-loop agent does not adapt its future actions based on the actual new state. Instead, it continues to follow a pre-planned sequence of actions, which are based on the simulated state during planning. For example, ToT-DFS and ToT-BFS (Yao et al., 2023a) produce a predefined sequence of actions based on a search through a static search tree or graph. This sequence is intended to be executed exactly as planned. Ideally, open-loop planning requires that the environment will remain as anticipated throughout the execution. Such environments satisfies the following assumptions: a) environments are static, b)(no unforeseen events or changes will affect the execution (closed-world assumption), and c) deterministic transitions. 2) Closed-Loop (Online): In contrast, after the closed-loop agent executes an action $a_{t}$, it observes the outcome in the real world (i.e., the resulting state) and can adjust its future action $a_{t+1}$ based on the new state $s_{t+1}$. The plan evolves dynamically as new information becomes available. MCTS-based frameworks, except for LATS (Zhou et al., 2024a), can be open-loop because it generates a sequence of actions based on simulations and a static model of the environment at a given state. However, the agent can operate in a closed-loop manner if MCTS is re-run at each new state: after a plan is generated for $s_{t}$, only $a_{t}$ is selected and executed. Instead of executing the rest $a_{t+1},\ldots$, the agent re-runs MCTS from the new state $s_{t+1}$ to determine the next best action $a_{t+1}$. These frameworks are suitable for task execution in dynamic and interactive environments.
LATS Requires Action Undoing
LATS (Zhou et al., 2024a) relies on direct environment interaction for path simulation. It assumes actions are reversible so that the agent can revisit earlier states if necessary. This limits its applicability to environments where undoing actions is viable.
6.3 Performance
MCTS May Degenerate in Early Stages
Chen et al. (2024b) observe that if all candidate steps receive equal (or zero) scores initially, MCTS lacks a clear basis for distinguishing among branches, potentially leading to suboptimal partial-plan selection. The performance is worse than iterative refinement (Madaan et al., 2023).
6.4 Efficiency
General Running Time
As noted by Chen et al. (2024b), tree search can be 10–20 times slower than iterative refinement, especially if the evaluator (LMPE) has less than 90% discrimination accuracy. High-accuracy LMPEs are essential to prune the search tree effectively, thereby reducing the number of iterations needed.
LMPP Expansion and Path Simulation with Memory
Finally, maintaining a memory of explored nodes can avoid repeated sampling and simulation. If a given state $s$ has already produced certain actions, those child nodes can be cached for subsequent expansions. Similarly, for simulation, previously simulated results can be stored for future simulation. By reusing these cached outcomes, the framework reduces redundant calls to LMPP or LMPT, thereby improving both inference cost.
Unnecessary LMPP Use in Some Cases
Some tasks possess a small, tractable action space (e.g., the Game of 24). In such scenarios, exhaustive action retrieval may be cheaper than performing multiple LMPP inferences. Designers must weigh the inference costs of LMPP against potential benefits, as LLM-based sampling can be expensive relative to enumerating a finite set of actions. The potential cost of the following LMPE+ evaluation should also be considered.
7 Related Work
Although the primary focus of this survey is on test-time compute via search, several related directions fall outside our current scope:
  • LLM Fine-Tuning for LMPRs. Recent methods adapt large language models through fine-tuning or preference optimization to enhance their policy, evaluation, or transition roles. Examples include Chain of Preference Optimization (Zhang et al., 2024), AlphaZero-like tree search (Wan et al., 2024), and AgentQ (Putta et al., 2024).
  • Search with Multi-Modal LLMs. Some work extends tree-based exploration and action selection to multi-modal contexts by incorporating visual features alongside textual reasoning steps, e.g., Mulberry (Yao et al., 2024)
  • Branching Without Search. Some frameworks utilize branching or tree-like expansions but do not incorporate a full-fledged search algorithm. Examples include Tree-Planner (Hu et al., 2024), Boost-of-Thoughts (Chen et al., 2024a), and Graph-of-Thoughts (GoT) (Besta et al., 2024). Although they adopt branching structures similar to traditional search, these methods rely on aggregation, sorting, or heuristics rather than explicit search procedures.